Consider the following production function.

TPL = 12L2 – 0.8L3

Determine the marginal product function(MPL)

Determine the average product function (APL)

Find the value of L that maximizes TPL

Find the value of L that maximizes APL

Find the value of L that maximizes MPL


1
Expert's answer
2021-09-07T09:45:38-0400

"TP_L=12L^2 -0.8L^3 \\

i) \; MP_L= \frac{∂(TP_L)}{∂(L)} \\

MP_L= 2 \times 12L^{(2-1)} – 3 \times 0.8L^{3-1} \\

MP_L = 24 L -2.4L^2 \\

ii) \; AP_L= \frac{TP_L}{L} \\

AP_L= \frac{12L^2 -0.8L^3}{L} \\

AP_L= 12L -0.8L^2"

iii) Maximozation of "TP_L"

Put "\frac{∂(TP_L)}{∂(L)} = 0" for maximum value

"24L-2.4L^2=0 \\

L(24-2.4L)=0 \\

24=2.4L \\

L= \frac{24}{2.4}=10 \;units"

iv) Put "\frac{∂(AP_L)}{∂(L)} =0" , for maximum value

"AP_L= 12L -0.8L^2 \\

\frac{∂(AP_L)}{∂(L)} = 12 -1.6L \\

12-1.6L=0 \\

L=\frac{12}{1.6}= 7.5 \;units"

v) Put "\frac{∂(MP_L)}{∂(L)}=0" for maximum value

"MP_L=24L -2.4L^2 \\

\frac{∂(MP_L)}{∂(L)} = 24-4.8L \\

24-4.8L=0 \\

L=\frac{24}{4.8}= 5 \;units"